Willie has an average (mean) of 85.6 on the first five exams. What is the least he can score in order to have an average of 85 for all six exams?
Answer by macston(5194)

The sum of scores on the first five exams:
(sum of scores)/(number of scores)=average
(sum of scores)/(5 exams)=85.6
sum of scores=428
.
To average 85 on 6 exams:
(sum of scores)/(6 exams)=85
sum of scores=510
.
Current sum=428
Desired sum=510
Needed score=Desired sum-current sum=510-428=82
.
ANSWER: Willie must score 82 to average 85 over all his exams.
.